OpenStudy (anonymous):

The space shuttle travels at 28000 mph while orbiting the earth. how far does the shuttle travel in feet in a day

OpenStudy (lgbasallote):

use dimension analysis thingy \[\LARGE \frac{28,000 \; \;\text{miles}}{\text{hour}} \times \frac{5,280 \;\;\text{feet}}{1 \; \;\text{mile}} \times \frac{24\;\; \text{hours}}{1 \;\; \text{day}}\] \[\LARGE\implies \frac{28,000 \; \;\cancel{\text{miles}}}{\cancel{\text{hour}}} \times \frac{5,280 \;\;\text{feet}}{1 \; \; \cancel{\text{mile}}} \times \frac{24\;\; \cancel{\text{hours}}}{1 \;\; \text{day}}\] does that help?

OpenStudy (anonymous):

\[28000*24*5280=3,548,160,000\text{ feet}\]